Consider the following payoff table.

S1

S2

S3

S4

A1

22

14

10

-8

A2

12

6

16

-2

A3

-4

8

10

12

With alpha = 0.4, the Hurwicz value for A3 would be:

4.8

2.4

10.4

5.6

With a maximin strategy, the best alternative would be:

A1

A2

A3

There is not enough information available.
Answer #1

1)Hurwicz value for A3 would be =max(of all values in A3)*0.4+min( all values in A3)*0.6

=12*0.4+(-4)*0.6=2.4

2) With a maximin strategy, the best alternative would be A2